Jane is three times the age of Pam,sue is two times the age of Jane,combined the ages equal 60, what is the age of each girl?
Answer by addingup(3677)

1/3J = P
2J = S
~~~~~~~~~~~~~~~~~~~~~~~~~~
J+P+S = 60
J+1/3J+2J = 60
3J+J+6J = 180
10J = 180
J = 18
Jane is 18
18/3 = 6
Pam is 6
18*2 = 36
Sue is 36
~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
18+6+36 = 60 Correct